# DaLPSR: Leverage Degradation-Aligned Language Prompt for Real-World Image Super-Resolution

### v1

@@ I Introduction
    We have proposed an effective and computational affordable one-stage degradation aligned multimodal framework for real-world image super-resolution task. Complementary priors including semantic content descriptions and degradation prompts are leveraged for accurate, fine-grained, and high-fidelity image restoration. Much richly tailored prompts for MLLMs elicit high-level semantic priors closely aligned with human perception.
@@
    We have conducted comprehensive experiments on popular benchmark datasets, including both synthetic and real-world cases. The experiment results according to reference-based and reference-free metrics have demonstrated that the proposed method achieves new state-of-the-art perceptual performance on real-world cases.
@@ III-B2 Image-Restoration Prompt Alignment Decoder
Specifically, in cross-attention layer, degradation prompts function as queries, and image features encoded from precede block serve as keys and values. During pretraining stage, we take the alignment process as fine-grained retrieval process. For each degradation type, we select the most relevant degradation prompt from the define four intervals.
Moreover, to avoid combination explosion problem, we respectively construct an image-restoration prompt alignment decoder for each degradation type. And then concatenate respective degradation prompt retrieved to form the final prompt description. Consequently, in this way, we can prioritize efficiency in computational resources, emphasize the role of degradation prompt and preserve good scalability to more degradation types in future.
@@ III-C Multi-Modality Semantic Priors Guidance
Benefiting from significantly scaling up the size and complexity of both data and language model, multimodal large language models (MLLMs) have achieved unprecedentedly amazing abilities on many high-level tasks, such as image captioning. In this work, we leverage a state-of-the-art MLLM known as LLaVA[50] to enhance the perception of LR images and extract informative semantic priors. LLaVA’s robust capabilities enable it to capture intricate relationships and patterns within visual data.
To enhance the coherence between the semantic content of LR images and the descriptions generated by LLaVA, we propose to refine the feature extraction ability of image encoder $\mathcal{E}$ on LR image. We conduct fine-tuning on image encoder $\mathcal{E}$ through minimizing the loss $\mathcal{L}_{\mathcal{E}}=\|\mathcal{E}(I_{\text{LR}})-\mathcal{E}(I_{\text{HR}})\|_{2}^{2}$, thereby reducing the disparity between LR and HR images.
As we know, it is essential to provide clear and comprehensive instructions to MLLM, ensuring that its generated outputs align with human intentions and preferences. Considering the robust capability of RAM[54] on image tag recognition, we propose to employ RAM to detect tags for LR images. Subsequently, we leverage the identified tags to formulate tailored instructions corresponding to different LR images. As depicted in Figure 4, the tags extracted from LR image serve as part of instructive prompts for LLaVA. Through the application of specific prompt instructions, LLaVA is capable of producing high-level semantic image descriptions that align with human perceptual understanding.
@@ IV-A2 Test Datasets
ICCV 2021 ICCV 2021 CVPR 2022 MM 2022 arxiv 2023 NeurIPS 2023 arxiv 2023 arxiv 2023 CVPR 2024 CVPR 2024
@@ IV-D1 Efficacy of Image-Restoration Prompt Alignment Decoder
As demonstrated in Table III, the Image-Restoration Prompt Alignment Decoder (IRPAD) significantly enhances the perceptual quality of restored LR images. This improvement is evident across various evaluation metrics. It underscores the decoder’s effectiveness in producing visually superior results.
@@
For more intuitive observation, we provide visual comparisons for ablation experiments. As illustrated in Figure 8, the implementation of IRPAD markedly diminishes image noise and mitigates aliasing artifacts. The resulted image displays textures with well-defined edges. The improvements explicitly demonstrate the decoder’s capability to produce visually appealing and high-fidelity restorations.
@@ IV-D2 Efficacy of Multi-Modality Semantic Priors
As the results presented in Table III, it is evident that the exclusion of high-level semantic priors results in a significant reduction in fidelity metrics. This decrease underscores the critical role that high-level priors play in preserving the perceptual quality and structural integrity of restored images.
Moreover, as illustrated in Figure 8, the effectiveness of utilizing multi-modality semantic priors is further validated. The incorporation of multi-modality semantic priors leads to images with clearer details and more refined textures, resulting in higher level of realism and overall image quality.
Both ablation studies have demonstrated the significant advantage of leveraging multi-modality prompts for superior image restoration outcomes.

### v2

@@ I Introduction
    We have proposed an effective and computational affordable one-stage degradation-aligned multimodal framework for real-world image super-resolution task. Complementary priors including semantic content descriptions and degradation prompts are leveraged for accurate, fine-grained, and high-fidelity image restoration. Much richly tailored prompts for MLLMs elicit high-level semantic priors closely aligned with human perception.
@@
    We have conducted comprehensive experiments on popular benchmark datasets, including both synthetic and real-world cases. The experiment results according to reference-based and reference-free metrics have demonstrated that the proposed method achieves a new state-of-the-art perceptual performance on real-world cases.
@@ III-B2 Image-Restoration Prompt Alignment Decoder
The alignment decoder takes degradation prompts and image features as inputs. In the cross-attention layer, these degradation prompts act as queries, while the image features serve as keys and values. During the pretraining stage, the process results in an output that is the degradation prompt most closely aligned with the input image, ensuring precise relevance and alignment.
Throughout the training process of the alignment decoder, the task is structured as a fine-grained retrieval problem. In the pretraining phase, the model is responsible for selecting the most pertinent degradation prompt for each degradation type from the predefined intervals. This involves learning to effectively align degradation prompts with image features, enabling the retrieval of the most relevant prompts based on the provided image features. Consequently, this approach allows us to optimize computational efficiency, highlight the significance of degradation prompts, and maintain scalability for accommodating additional types of degradation in the future.
@@ III-C Multi-Modality Semantic Priors Guidance
By significantly scaling up the size and complexity of both data and language models, multimodal large language models (MLLMs) have achieved remarkable performance in high-level tasks like image captioning. In our work, we utilize a cutting-edge MLLM called LLaVA [50] to enhance the perception of low-resolution (LR) images and extract valuable semantic information. LLaVA’s advanced capabilities allow it to capture complex relationships and patterns within visual data, making it exceptionally effective for this purpose.
To improve the alignment between the semantic content of LR images and the descriptions generated by LLaVA, we propose enhancing the feature extraction capability of the image encoder $\mathcal{E}$ for LR images. This is achieved by fine-tuning the image encoder $\mathcal{E}$ to minimize the loss function $\mathcal{L}{\mathcal{E}}=|\mathcal{E}(I{\text{LR}})-\mathcal{E}(I_{\text{HR}})|_{2}^{2}$, thereby reducing the differences between LR and high-resolution (HR) images. This process ensures that the encoder extracts features from LR images that are closer to those of HR images, leading to more accurate semantic interpretations.
Furthermore, providing clear and comprehensive instructions to MLLMs is crucial to ensuring that their outputs align with human intentions and preferences. Given the powerful capabilities of RAM [54] in image tag recognition, we employ RAM to detect relevant tags for LR images. These identified tags are then used to create specific instructions tailored to each LR image. As illustrated in Figure 4, the tags extracted from an LR image become part of the instructive prompts for LLaVA. By utilizing these targeted prompt instructions, LLaVA can produce high-level semantic image descriptions that resonate with human perceptual understanding, thereby enhancing the coherence and relevance of the generated content.
@@ IV-A2 Test Datasets
ICCV 2021 ICCV 2021 CVPR 2022 MM 2022 IJCV2024 NeurIPS 2023 ECCV 2024 arxiv 2023 CVPR 2024 CVPR 2024
@@ IV-D1 Efficacy of Image-Restoration Prompt Alignment Decoder
As shown in Table III, the Image-Restoration Prompt Alignment Decoder (IRPAD) greatly improves the perceptual quality of restored low-resolution (LR) images, with enhancements evident across various evaluation metrics. This underscores the decoder’s effectiveness in producing visually superior results. To provide a clearer understanding, we include visual comparisons from our ablation experiments. As depicted in Figure 8, the use of IRPAD significantly reduces image noise and minimizes aliasing artifacts, resulting in textures with well-defined edges. These improvements clearly demonstrate the decoder’s capability to achieve visually appealing and high-fidelity restorations.
@@ IV-D2 Efficacy of Multi-Modality Semantic Priors
As shown in Table III, excluding high-level semantic priors results in a noticeable drop in fidelity metrics. This decrease emphasizes the essential role these priors play in maintaining the perceptual quality and structural integrity of restored images. Furthermore, as depicted in Figure 8, the effectiveness of using multi-modality semantic priors is clearly validated. Incorporating these priors leads to images with sharper details and more refined textures, enhancing realism and overall image quality.
In the ablation study, to ensure a fair comparison, we constructed baseline models by systematically excluding IRPAD and/or MLLM components. For the baseline without IRPAD, we removed the IRPAD module and relied on a basic image restoration pipeline involving standard denoising, lacking the refined adjustments provided by IRPAD. For the baseline without MLLM, we excluded MLLM components, processing images based solely on intrinsic features without multimodal inputs or semantic prompts, using only basic image characteristics for restoration. These baselines were trained and tested under identical conditions, using the same datasets and evaluation metrics such as LPIPS and CLIPIQA, to ensure comparability. By analyzing these models, we clearly demonstrate the specific contributions of IRPAD and MLLM to the overall restoration quality.
@@ V Conclusion
Acknowledgment
This work is supported by National Natural Science Foundation of China under Grand No. 62366021, and Jiangxi Provincial Graduate Innovation Funding Project under Grand YC2023-S298.